OpenStudy (anonymous):

The table shows the height of a plant as it grows. What equation in point-slope form gives the plant’s height at any time? Let y stand for the height of the plant in cm and let x stand for the time in months. Time (months): 3; Plant Height (cm): 15 Time (months): 5; Plant Height (cm): 25 Time (months): 7; Plant Height (cm): 35 Time (months): 9; Plant Height (cm): 45

OpenStudy (anonymous):

A. y – 15 = (x – 3) B. y – 15 = 5(x – 3) C. y – 3 = (x – 15) D. The relationship cannot be modeled. A, B, C, or D? Explain your answer. (:

OpenStudy (nubeer):

A.. Pick any two lines and calculate the gradient. i will pick first 2 for it m = (25-15)/(5-3) m = 5 y-y1 = m(x-x1) y -15 = 5( x-3)

OpenStudy (anonymous):

Thank you for your answer. (:

OpenStudy (nubeer):

you're welcome :)